40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7440     Entry time: Wed Sep 26 01:10:34 2012
Author: Jenne 
Type: Update 
Category: IOO 
Subject: PZT2 not working?!?! MC back to normal 

[Jenne, Evan, Den]

MC REFL beam is back on the PD, and the mode cleaner locks.  It looks like we're a little high on the MC Refl camera, but the MC spots were measured, and don't look like they changed from Friday (or maybe Monday?), the last time they were measured. We took this to be acceptable MC alignment, and did not touch the PSL table's pointing.

The laser power reduction optics were removed, and placed out of the way on the PSL table (where do they belong?).  PSL-POS and PSL-ANG aren't quite perfectly centered, but a beam dump had been in the way of that path, so I don't know if they drifted bad, or if it was a sudden thing.  The beam is still hitting the QPDs though.  After removing the beam power reducing optics, we recentered the MC REFL beam on the REFL PD, still not touching any PSL alignment.  MC mirrors were aligned (Den did this work while I showed Evan around, so I don't know by how much), and MC Trans was maximized (really MC Refl was minimized, making sure that the unlocked MC Refl was the usual 4.something units on the EPICS readback.

We turned on the PZT high voltage supplies for the output steering PZTs and for the input steering PZTs.  We left the OMC locking PZT supplies off, since we're still not using the OMC.  Sadly, the beam coming out of the AS port looks clipped somewhere.  [SELF: attach the videocapture shot when you get to work tomorrow] We tried moving PZT2's sliders, but nothing happened!!! I can move BS and the ITMs to get the beam mostly unclipped, but I really need to be able to move the PZTs, or at least one of them.  IPPOS and IPANG beams are hitting the QPDs (although they're not centered perfectly), so the PZTs came back mostly to the same positions, but not exactly.  Evan and I sat next to the input steering PZT controllers in 1Y3, and moved the sliders around.  For most of the range, nothing changes on the LCD screen for either PZT2 pitch or yaw.  Yaw can make 2 small steps near the far negative side of the slider, but nothing happens for most of the slider.  Pitch really doesn't do anything for any part of the slider.  We ensured that the LED labeled "CL ON" was not illuminated, next to the button labeled "closed loop", for all 4 controllers (PZT1 and 2, pitch and yaw).  Sad!!   I don't know if the LCD screen on the front panel of the PZT controllers is a readback of signal supplied to the PZTs, or of the strain gauges.  I really hope it's the controller that's not working, rather than the PZTs themselves.  The PZTs were fine before we vented, and Koji and I did our centering of the PZT range check during the vent, so they were fine then.  What happened???  All PZT high voltage supplies were off during the pump-down.  I turned them off yesterday, and Evan and I turned them back on tonight around 9:30pm or 10pm.  What else could make them bad?

Without being able to move PZT2, just using BS and / or ITMs, I was unable to completely make the beam look nice on the AS camera.  I came close, but it still seems a little bit funny, and I had to move the suspended optics quite a bit to find that place.  This is not good. 

ELOG V3.1.3-